OpenStudy (ryguy):

please help! Find three consecutive odd integers such that the sum of the largest and twice the smallest is 25. If x represents the smallest integer, then which equation could be used to solve the problem?

OpenStudy (anonymous):

hey u have choices or u have to explain

OpenStudy (ryguy):

choices are x + 3 x + 4 x + 6

OpenStudy (ryguy):

wait those choices are the wrong ones sorry 2x + x + 2 = 25 2x + x + 4 = 25 2x + 2x + 4 = 25

OpenStudy (anonymous):

alright so u want an odd number

OpenStudy (anonymous):

Oh okay np

OpenStudy (ryguy):

three of them. i dont need to solve for x i just need to figure our which equation to use.

OpenStudy (anonymous):

u want one of the X's all alone in your equation

OpenStudy (anonymous):

and since you have x+2 and x+4 which answer choice has 2 and 4

OpenStudy (ryguy):

nvm got it